### v2

# A Mixture-of-Experts Framework with Log-Logistic Components for Survival Analysis on Histopathology Images

Abstract
We propose a modular framework for predicting cancer-specific survival from whole-slide pathology images (WSIs). The method integrates four components:(i) Quantile-Gated Patch Selection via quantile-based thresholding to isolate prognostically informative tissue regions; (ii) Graph-Guided Clustering using $k$-NN graph to capture phenotype-level heterogeneity through spatial–morphological coherence; (iii) Hierarchical Context Attention to learn intra- and inter-cluster interactions; and (iv) an Expert-Driven Mixture of Log-logistics framework to estimate complex survival distributions using Log-logistics distributions. The model attains a concordance index of $0.644{\pm}0.059$ on TCGA-LUAD, $0.751{\pm}0.037$ on TCGA-KIRC, and $0.752{\pm}0.011$ on TCGA-BRCA, respectively, outperforming existing state-of-the-art approaches.
Index Terms—  Survival Analysis, Whole-Slide Images, Histopatholgy, Attention, Mixture of Experts, Deep Learning.
1 Introduction
Accurate survival prediction from histopathology images is critical for personalized cancer care, guiding treatment and risk stratification. Whole-slide images (WSIs) capture rich tissue-level morphology—including tumor architecture and microenvironmental context—but their massive size and lack of localized annotations make direct modeling challenging. Patch-based multiple instance learning (MIL) has thus become the dominant paradigm, dividing each slide into smaller tiles for weakly supervised analysis.
While MIL has succeeded in cancer grading and subtyping, survival prediction requires modeling long-range dependencies and subtle morphological cues across spatially distributed tissue regions. Conventional statistical approaches such as the Cox proportional hazards model cannot capture these non-linear, high-dimensional relationships, whereas deep models often lack interpretability and spatial reasoning.
To address these challenges, we present a unified framework combining four modules: (1) Quantile-Gated Patch Selection dynamically identifies patches and retains prognostically relevant tissue regions through quantile-based filtering. (2) Graph-Guided Clustering of Relevant Patches to capture spatial morphological coherence; (3) Hierarchical Context Attention to learn both intra- and inter-cluster dependencies; and (4) expert-driven mixture of Log-logistics framework for survival distribution estimation. Together, these components enable interpretable and accurate survival prediction by jointly leveraging spatial structure, phenotype abstraction, and probabilistic outcome modeling.
2 Related Work
Weakly supervised learning has been central to WSI-based survival modeling, primarily through multiple instance learning (MIL). Early models such as ABMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], and DSMIL [li2021dualstreammultipleinstancelearning] used attention pooling to highlight prognostic patches but overlooked spatial context. Graph-based PatchGCN [chen2021slideimages2dpoint] captured spatial adjacency, while hierarchical transformers like HGT [10.1007/978-3-031-43987-2_72], TransMIL [shao2021transmiltransformerbasedcorrelated], and HIPT [chen2022scalingvisiontransformersgigapixel] modeled multi-scale features at high computational cost. Recent multimodal methods such as PathoGen-X [10981028], CoC [ZhoHai_CoC_MICCAI2025], and HiLa [CuiJia_HiLa_MICCAI2025] integrate histopathology with genomic, methylation, and language data, while SCMIL [Yang_2024] estimates Gaussian distribution-based survival predictions. However, these approaches lack mechanisms to disentangle latent phenotypes or exploit structural coherence. Our expert-driven mixture of log-logistics module addresses this gap by integrating phenotype-aware clustering with expert-specific survival estimation, where specialized experts capture distinct features and a gating network adaptively fuses their outputs for improved interpretability.
Fig. 1: (a) Proposed survival framework. (b) Expert-guided Mixture Density Modeling block.
3 Methodology
[…]
$\mathbf{P}_{\text{feat}}\in\mathrm{R}^{n\times d},$ (1)
where $n$ denotes the number of extracted patches and $d$ the feature dimensionality.
The proposed framework comprises four modules: Quantile-Gated Patch Selection(QGPS), Graph-Guided Clustering(GGC), Hierarchical Context Attention(HCA), and Expert-Driven Mixture of Log-logistics Module(EDMLL).
3.1 Quantile-Gated Patch Selection (QGPS)
To focus on diagnostically informative regions, a quantile-based learnable Multi Layer Perceptron(MLP) selects task-relevant patches according to their importance scores derived from model logits. For each WSI, a threshold $\tau_{q}$ is determined as the $q$-quantile of logits, retaining the top $(1-q)\times 100\%$ of patches:
$\mathcal{P}_{\text{sel}}=\{\mathbf{P}_{i}\mid\text{logit}_{i}>\tau_{q}\},\quad\mathcal{P}_{\text{rem}}=\{\mathbf{P}_{i}\mid\text{logit}_{i}\leq\tau_{q}\}.$ (2)
This adaptive selection removes noisy or background patches, ensuring downstream modules focus on semantically meaningful regions relevant to survival outcomes.
3.2 Graph-Guided Clustering (GGC)
The selected patches $\mathcal{P}_{\text{sel}}\in\mathrm{R}^{m\times d}$ are clustered into morphologically and spatially coherent groups. First, patch features and coordinates are normalized to comparable scales. Morphological similarity between patches is computed using cosine similarity:
$S_{\text{morph}}(i,j)=\frac{\langle\mathbf{P}_{i},\mathbf{P}_{j}\rangle}{\|\mathbf{P}_{i}\|\,\|\mathbf{P}_{j}\|},$ (3)
while spatial similarity is defined via an exponential kernel over Euclidean distances $D_{ij}$ between normalized coordinates:
$S_{\text{spatial}}(i,j)=\exp\!\left(-\frac{D_{ij}}{\sigma_{D}}\right),\quad\text{where }\sigma_{D}=\text{std}(D)+\varepsilon.$ (4)
A composite similarity matrix balances both cues:
$S=\omega_{\text{morph}}S_{\text{morph}}+\omega_{\text{spatial}}S_{\text{spatial}}$ (5)
Based on $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and patches are then clustered into $G$ groups $L_{1},\dots,L_{G}$} using GPU-accelerated K-Means, guided by the graph structure. Each cluster represents a local tissue microenvironment such as tumor epithelium, stroma, or necrosis, promoting a biologically meaningful and spatially coherent organization across the slide.
3.3 Hierarchical Context Attention (HCA)
To capture relationships within clutsers, a two-level attention mechanism is adopted.
Intra-cluster attention. Each cluster $L_{i}$ undergoes multi-head self-attention (MHSA) [vaswani2017attention] to model fine-grained local dependencies among patches:
$L^{\prime}_{i}=\text{LayerNorm}\!\big(L_{i}+\text{MHSA}(L_{i})\big).$ (6)
This step enhances contextual awareness within morphologically similar regions.
Inter-cluster attention. The refined clusters are summarized into representative embeddings
$R_{i}=\frac{1}{|L^{\prime}_{i}|}\sum_{\mathbf{x}\in L^{\prime}_{i}}\mathbf{x},$ (7)
which are then processed by another MHSA layer to capture global relationships across regions:
$R^{\prime}=\text{LayerNorm}\!\big(R+\text{MHSA}(R)\big).$ (8)
Intra-cluster representations are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{g})_{g=1}^{G}$. A broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, derived by averaging $R^{\prime}$ across clusters and expanding to match $\widetilde{P}$, is added residually to form $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with remaining irrelevant patch features to yield $\mathcal{P}_{\text{final}}$.
WSI-level aggregation. Slide-level representation is derived via attention pooling as in AMIL [ilse2018attentionbaseddeepmultipleinstance]
$\mathbf{z}_{\text{WSI}}$ $=\sum_{i}\alpha_{i}\,\mathcal{P}_{\text{final},i},$ (9)
$\alpha_{i}$ $=\text{softmax}\!\big(W_{a}\,\tanh(W_{h}\,\mathcal{P}_{\text{final},i}^{\top})\big),$
where $\mathbf{z}_{\text{WSI}}$ represents the integrated prognostic feature for the WSI.
3.4 Expert-Driven Mixture of Log-logistics
An Expert-Driven Mixture of Log-Logistics module (Figure 1b) models survival probabilities, with each expert specializing in distinct WSI features $\mathbf{z}_{\text{WSI}}$. Built on the mixture-of-experts framework [bishop2006pattern], it uses multiple experts to form Log-logistics Distributions(LLD) that capture heterogeneous survival patterns. The log-logistic probability density is given by
$\mathrm{LLD}(t\mid\alpha,\beta)=\frac{(\beta/\alpha)(t/\alpha)^{\beta-1}}{\big[1+(t/\alpha)^{\beta}\big]^{2}},\quad t>0,$ (10)
where $\alpha,\beta>0$ denote the scale and shape parameters.
The survival time $t$ of a whole slide image is directly modeled on its features $\mathbf{z}_{\text{WSI}}$ derived via the attention pooling layer, as shown in Fig. 1(b). Each expert $e$ in the mixture models the conditional distribution over $t$ using a log-logistic mixture model (LLD):
$p(t\mid\mathbf{z}_{\text{WSI}},e)=\sum_{k=1}^{K}\lambda^{(e)}_{k}(\mathbf{z}_{\text{WSI}})\,\mathrm{LLD}\!\big(t\mid\alpha^{(e)}_{k},\beta^{(e)}_{k}\big)$, (11)
where each $k$ component follows a log-logistic distribution with scale $\alpha^{(e)}_{k}$ and shape $\beta^{(e)}_{k}$, while mixture weights $\lambda^{(e)}_{k}$ are computed using a softmax over the $K$ components. Learnable global anchors $(P_{\alpha},P_{\beta})\in\mathrm{R}^{K}$, corresponding to scale and shape parameters, are projected for each expert as below:
$\alpha^{(e)}=\text{softplus}\!\big(W^{(e)}_{\alpha}P_{\alpha}\big),\quad\beta^{(e)}=\text{softplus}\!\big(W^{(e)}_{\beta}P_{\beta}\big)$ (12)
A gating network $G(\mathbf{z}_{\text{WSI}})$ assigns softmax weights to experts, producing the overall time probability density function (TPDF) and survival probability function (SPF) as:
$\text{TPDF}(t\mid\mathbf{z}_{\text{WSI}})=\sum_{e,k}G_{e}(\mathbf{z}_{\text{WSI}})\lambda^{(e)}_{k}\,\mathrm{LLD}\!\big(t\mid\alpha^{(e)}_{k},\beta^{(e)}_{k}\big)$, (13)
$\text{SPF}(t\mid\mathbf{z}_{\text{WSI}})=1-\!\sum_{e,k}G_{e}(\mathbf{z}_{\text{WSI}})\lambda^{(e)}_{k}F_{\mathrm{LLD}}\!\big(t\mid\alpha^{(e)}_{k},\beta^{(e)}_{k}\big)$. (14)
where $F_{\mathrm{LLD}}$ is obtained by integrating the LLD.
3.5 Training Objective and Regularization
The model is trained using a negative log-likelihood loss ($\mathcal{L}_{\text{NLL}}$) that handles both censored ($c=0$ i.e. event not yet occurred) and uncensored ($c=1$ i.e. event occurred) data. For an observed time $t_{d}$ and censoring indicator $c\in\{0,1\}$,
$\mathcal{L}_{\text{NLL}}=-\,c\,\log\!\big(\text{TPDF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)-(1-c)\,\log\!\big(\text{SPF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)$
To promote regularization, a gating entropy loss is introduced($\mathcal{L}_{\text{ent}}$).
$\mathcal{L}_{\text{ent}}=-{\textstyle\sum_{e}}G_{e}(\mathbf{z}_{\text{WSI}})\log\!\big(G_{e}(\mathbf{z}_{\text{WSI}})\big).$ (16)
The overall loss combines the two above terms:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{NLL}}+\lambda_{\text{ent}}\,\mathcal{L}_{\text{ent}}.$ (17)
Fig. 2: Kaplan–Meier survival curves showing significant high- vs. low-risk separation across TCGA cohorts.
Table 1: Left: TCGA performance (TDC, higher is better). Right: Ablations on $q$ in QGPS, GGC, HCA and $e$ in EDMLL. Best TDC results in bold.
(a) TCGA results. Modality:I=image,G=genomic,M=methylation,L=language prompt.
Method Modality LUAD KIRC BRCA MEAN
AMIL [ilse2018attentionbaseddeepmultipleinstance] I $0.632\pm 0.021$ $0.691\pm 0.022$ $0.735\pm 0.013$ $0.686$
CLAM [lu2020dataefficientweaklysupervised] I $0.615\pm 0.041$ $0.680\pm 0.028$ $0.652\pm 0.027$ $0.649$
DSMIL [li2021dualstreammultipleinstancelearning] I $0.609\pm 0.039$ $0.644\pm 0.013$ $0.668\pm 0.011$ $0.640$
TransMIL [shao2021transmiltransformerbasedcorrelated] I $0.568\pm 0.017$ $0.642\pm 0.076$ $0.636\pm 0.018$ $0.615$
PatchGCN [chen2021slideimages2dpoint] I $0.587\pm 0.019$ $0.675\pm 0.045$ $0.582\pm 0.037$ $0.615$
HIPT [chen2022scalingvisiontransformersgigapixel] I $0.549\pm 0.025$ $0.640\pm 0.037$ $0.625\pm 0.046$ $0.605$
HGT [10.1007/978-3-031-43987-2_72] I $0.607\pm 0.058$ $0.648\pm 0.018$ $0.648\pm 0.022$ $0.634$
SCMIL [Yang_2024] I $0.622\pm 0.015$ $0.688\pm 0.037$ $0.674\pm 0.048$ $0.661$
OTSurv [RenQin_OTSurv_MICCAI2025] I $0.638\pm 0.077$ $0.750\pm 0.149$ $0.621\pm 0.071$ $0.670$
PathoGen-X [10981028] I+G $0.620\pm 0.008$ --- $0.670\pm 0.020$ $0.645$
CoC [ZhoHai_CoC_MICCAI2025] I+M --- $0.709\pm 0.048$ $0.654\pm 0.036$ $0.681$
HiLa [CuiJia_HiLa_MICCAI2025] I+L $0.643\pm 0.055$ --- $0.659\pm 0.044$ $0.651$
Ours I 0.644 $\pm$ 0.059 0.751 $\pm$ 0.037 0.752 $\pm$ 0.011 0.716
(b) Ablations.
Variant LUAD KIRC BRCA
[…]
4 Experimental Setup
4.1 Datasets
We evaluate the framework on three TCGA cohorts: Lung Adenocarcinoma (LUAD; 459 WSIs), Kidney Renal Clear Cell Carcinoma (KIRC; 509 WSIs), and Breast Invasive Carcinoma (BRCA; 956 WSIs) [tcga2013pan]. Each dataset contains WSIs annotated with overall survival time, measured in months from diagnosis to death (uncensored data, c = 1) or last follow-up (censored data, c = 0).
4.2 Training Configuration
The quantile hyperparameter is set to $q=0.25$, retaining the top $75\%$ of patches based on logits. The $k$-NN graph constructed from the similarity matrix $S$ connects each point to its 10 nearest neighbors. Based on the graph connectivity, the patch indices are clustered into $G$ groups of 64 each, and these grouped patches are subsequently processed by an 8-head MHSA within the HCA module. The Expert-Driven Mixture of Log-Logistics module employs five experts, each modeling a log-logistic distribution (LLD) with $K=100$ components. Model is trained for 20 epochs with Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) and a batch size of 1. Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric over 5-fold cross-validation. All experiments are conducted on an NVIDIA A6000 GPU.
5 Results and Analysis
5.1 Quantitative Results
Table 1(a) summarizes performance across TCGA datasets. Our framework consistently achieves the highest mean TDC, indicating superior discrimination and calibration. Histology-based multiple instance learning (MIL) models—AMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], DSMIL [li2021dualstreammultipleinstancelearning], and TransMIL [shao2021transmiltransformerbasedcorrelated] served as backbones for the expert-driven mixture of log-logistics block, with performances reported for comparison. Survival-specific MIL methods such as HIPT [chen2022scalingvisiontransformersgigapixel], Patch-GCN [10981028], HGT [10.1007/978-3-031-43987-2_72], SCMIL [Yang_2024] and OTSurv [RenQin_OTSurv_MICCAI2025] act as baselines. Since multimodal frameworks [10981028, ZhoHai_CoC_MICCAI2025, CuiJia_HiLa_MICCAI2025] lack public code, their results are taken from the respective papers. By combining quantile-based patch selection, graph-guided clustering, hierarchical attention, and expert-driven mixture of log-logistics framework—each expert capturing distinct morphological or contextual cues—our model effectively identifies survival-relevant regions and achieves state-of-the-art performance on all datasets. Remarkably, it even surpasses multimodal systems while relying solely on histopathology images.
5.2 Ablation Studies and Interpretability
Ablation studies are conducted to evaluate the contributions of Quantile-Gated Patch Selection, Graph-Guided Clustering, Hierarchical Context Attention and Expert-Driven Mixture of Log-logistics(Table 1(b)). Varying the quantile parameter showed that q=0.25q{=}0.25 provided the best trade-off, preserving prognostic regions while filtering irrelevant tissue. Eliminating any module resulted in a reduction in TDC, confirming their contributions to noise suppression and contextual feature integration. Single-expert model fell short, whereas five experts yielded superior results, underscoring the benefit of multi-expert modeling. For interpretability, patients were categorized into high and low-risk groups based on their predicted survival scores. The Kaplan–Meier curves(Fig. 2) revealed distinct survival trends with significant separation (log-rank p=0.046p{=}0.046 for LUAD, p=0.005p{=}0.005 for KIRC, p=0.017p{=}0.017 for BRCA), confirming that the expert-driven architecture effectively captures prognostically relevant morphological patterns.
6 Conclusion
We proposed a unified framework for WSI-based survival prediction integrating quantile-based patch selection, graph-guided clustering, hierarchical context attention, and expert-driven mixture of log-logistics. By modeling local–global tissue interactions through clustered attention and estimating survival distributions via multiple experts, the method improves calibration and discrimination. Experiments on TCGA-LUAD, TCGA-KIRC, and TCGA-BRCA show consistent gains in time-dependent concordance over pathology-based and multimodal baselines. Future work will extend the framework towards multimodal and uncertainty-aware patient-wise survival prediction for clinical reliability.
7 Compliance with ethical standards
This study used publicly available human data from [tcga2013pan]; hence, no ethical approval was required under the associated open-access license.
8 Acknowledgments
The results of this study are based on the data collected from the public TCGA Research Network [tcga2013pan].